OpenStudy (anonymous):

The two numbers taht are 5 units from 3 on the numberline are represented by the equation |n-3|=5. What are those two numbers?

OpenStudy (anonymous):

Well I will explain it to you, n-3=5 and n-3=-5 so what minus 3 equals 5, and what minus 3 equals negative 5?

OpenStudy (anonymous):

So you put 5+3=n and -5+3=n so the first one is 8 and the second one is -2

OpenStudy (anonymous):

You got it?

OpenStudy (anonymous):

yeah thanx